# Honeybee: Decentralized Peer Sampling with Verifiable Random Walks for Blockchain Data Sharding

### v1

@@ Abstract.
Data sharding—in which block data is sharded without sharding compute—is at the present the favored approach for scaling Ethereum. A key challenge toward implementing data sharding is verifying whether the entirety of a block’s data is available in the network (across its shards). A central technique proposed to conduct this verification uses erasure coded blocks and is called data availability sampling (DAS). While the high-level protocol details of DAS has been well discussed in the community, discussions around how such a protocol will be implemented at the peer-to-peer layer are lacking. We identify random sampling of nodes as a fundamental primitive necessary to carry out DAS and present Honeybee, a decentralized algorithm for sampling node that uses verifiable random walks. Honeybee is secure against attacks even in the presence of a large number of Byzantine nodes (e.g., 50% of the network). We evaluate Honeybee through experiments and show that the quality of sampling achieved by Honeybee is significantly better compared to the state-of-the-art. Our proposed algorithm has implications for DAS functions in both full nodes and light nodes.
blockchain, peer-to-peer, data availability sampling
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Symposium on Principles of Distributed Computing; June 17–21, 2024; Nantes, Franceprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Networks Network algorithmsccs: Theory of computation Distributed algorithmsccs: Networks Peer-to-peer protocolsccs: Networks Network performance evaluationccs: Networks Network propertiesccs: Networks Peer-to-peer networksccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Blockchains are steadily maturing into ecosystems that support decentralized applications (dapp) in diverse domains—including finance, payments, storage, games, healthcare etc.—and used by millions of clients conducting billions of dollars of transactions each day (CoinMarketCap, 2024). As demand for dapp usage increases, it is important that the blockchains can handle the high rate of transaction requests. Today the Ethereum blockchain can process at most a few ten transactions per second on average which has resulted in unacceptably high transaction fees during periods of high demand (Etherscan, 2024).
To address the scaling problem, the Ethereum community is focused on the development of a data-sharding design in which (1) block sizes are increased from the current 80 kB (average) to as big as 30 MB (Nikolaenko and Boneh, 2024) and erasure coded, (2) each validator node stores only a small chunk of each coded block. An increased block size naturally admits a greater number of transactions and improves transaction throughput of the chain. Combined with the use of layer-2 scaling methods, particularly Rollups (Ethereum.org, 2024e), Ethereum envisions a roadmap wherein a theoretical maximum throughput of 100,000 transactions per second is feasible (Forum, 2024). Data sharding in Ethereum is colloquially referred as Dank sharding, named after its proposer (EIP-4844, 2024).
@@
A fundamental challenge toward realizing data sharding is the design of the peer-to-peer (p2p) network. Unlike Ethereum 1.0 which used a simple broadcast gossip primitive to disseminate transactions and blocks in the network or Ethereum 2.0 which uses a publish-subscribe model to disseminate different types of messages (e.g., attestations, blocks, transactions) over different subnets, a node in a data-sharded Ethereum must additionally be able to check the availability of a full block by sampling random chunks from other nodes (Al-Bassam et al., 2018). This requires a requesting node to be able to contact randomly selected nodes in the network and download chunks. Having the ability to sample random chunks is important not only for full functional nodes, but also for light nodes (e.g., a wallet running on a smart phone) which can number in the millions compared to the thousands of full nodes available today (Ethernodes.org, 2024).
@@
In an open, permissionless and decentralized setting that Ethereum operates on, developing an algorithm by which a node can uniformly sample other nodes in the p2p network is far from straightforward. At present Ethereum’s p2p stack (based on devp2p and libp2p) use the Kademlia distributed hash table (DHT) protocol (Maymounkov and Mazieres, 2002) to sample and discover new nodes, as part of the node discovery procedure (to-peer Networking Specifications, 2024b, a). To discover a random node in the network, all a node has to do is issue a query for a randomly selected target identifier in Kademlia and receive the IP address information of the node that is closest to the target identifier (in Kademlia’s XOR distance sense).
@@
Even though Ethereum uses Kademlia as the protocol of choice to perform network-level node sampling, in this paper we argue that in the presence of adversarial nodes (e.g., Sybils), Kademlia does not achieve uniform sampling and can cause a requesting node to become eclipsed. As a potential replacement to Kademlia, in this paper we propose Honeybee which is a fully decentralized p2p algorithm for performing uniform node sampling even in the presence of adversarial nodes. A Honeybee node achieves sampling by participating in several random walks over the p2p overlay. Each node in Honeybee maintains an address table containing addresses of the most recently sampled peers which are also used to progress random walks of other peers visiting the node. To protect against adversarial attacks, Honeybee uses verifiable randomness derived from the blockchain to perform the walks. Additionally, Honeybee nodes also perform peer-to-peer reconciliation of node address tables to identify and expose attackers engaging in equivocating their address tables.
@@
We do not attempt to present a full-fledged p2p network design for data sharding in Ethereum in this paper. Rather we posit that a uniform node sampling capability will have a central role to play in the (eventual) overall network design. The sampling capability can also enhance the effectiveness of broadcast in today’s Ethereum p2p subnets. [^1] In the context of Ethereum’s publish-subscribe p2p network, a subnet is an overlay wherein all nodes are interested in the same topic and gossip messages on that topic. The emerging trend of modular blockchains advocates for a separate data availability layer in the blockchain stack, the implementation of which can make use of Honeybee as well (Celestia, 2024; Docs, 2024a).
@@
We evaluate Honeybee through a custom simulator we have built. [^2] Code will be made open source. Compared to the baseline algorithms GossipSub and Kademlia, Honeybee achieves the same level of near-uniform sampling when all nodes in the network adhere to the protocols. However, when the network contains adversarial nodes, Honeybee outperforms GossipSub and Kademlia by 4-63% in terms of sampling adversarial peer ratio. We define that an algorithm achieves $\epsilon$-uniform sampling when the sampling adversarial peer ratio from the algorithm is bounded from above by the sum of the true adversarial nodes ratio in the network and $\epsilon$. Under such standard, Honeybee consistently achieves 0.03-uniform sampling with 5-50% adversarial nodes in the network.
2. Background
2.1. Data availability sampling
Transactions per second, or throughput, is a key measure of blockchain scalability. Today the Ethereum blockchain has an average block size of less than 200 kB corresponding to a throughput of around 30 transactions per second. Increasing this throughput to 1000s of transactions per second (to the scale of Visa, for example) has been a long-standing open challenge in the community.
While several layer-2 solutions such as state channels, side channels, Plasma etc. (Ethereum.org, 2024b) have been proposed to improve throughput, a particularly important layer-2 solution that has become the focus of Ethereum’s scaling roadmap is the rollup (Ethereum.org, 2024e). Briefly, a rollup is an independent blockchain that is bridged to the Ethereum’s main chain via a smart contract. For increased security in the rollup chain, blocks produced in the rollup are periodically published on the main chain in a compressed form. This allows any verifier on the main chain to verify the rollup operation, and slash rollups block producers in case of a mistake. The rollups blocks published on Ethereum are stored as data ‘blobs’ and are not executed by Ethereum validators by default.
Including blobs from rollups within an Ethereum block can significantly amplify the size of the block to several megabytes. In this scenario, requiring all the nodes to store all the blocks (as is the case today) can overwhelm resource-limited nodes and affect the decentralization of the network. Therefore, Ethereum envisions a sharded design in which a block is split into many smaller chunks, and each node is required to store only a small number of the chunks (Ethereum.org, 2024a; Al-Bassam et al., 2018). The block header is stored by all the nodes.
The challenge is to ensure that a published block can be successfully reconstructed from its chunks later if required. E.g., a malicious block publisher may reveal only the block header, retaining the block body; or the publisher may publish only 90% of the chunks; or some of chunks can be modified during publication etc. To solve this, a block is first erasure coded to increase its size by 4 times (Ethereum uses a 2D-Reed Solomon code). The erasure coded block is then divided into chunks which are then dispersed over the validator network.
Erasure coding the blocks is advantageous in two ways:
  1. (1)
    It is possible for a node to reconstruct the entire block by fetching any random 25% of the chunks. This is easier to achieve compared to retrieving all 100% of the chunks if a block is not erasure coded.
  2. (2)
    It is simple for a node to verify that a block’s data is available (i.e., has been dispersed correctly). All the node needs to do is sample a small $k$ number of random chunks from the network. If the node can successfully retrieve all $k$ chunks, then with probability $>1-1/4^{k}$, block data is available over the network. This is because if a malicious node attempts to withhold publishing of a portion of the block, then it can do it only by withholding more than 75% of the coded chunks.
Correctness of each individual chunk is verified using polynomial commitments (specifically, KZG commitment) that are packaged along with the chunks. The process of checking whether adequate block data is present in the network by sampling is called data availability sampling.
Data availability sampling is essential not only for full nodes but is also beneficial to light nodes (Ethereum.org, 2024c). Light nodes are end users interacting with the blockchain through lightweight application software than can run easily on resource-limited devices. A simple example is a user running a wallet software on her browser that interacts with the blockchain through the wallet provider’s servers. As illustrated by this example, today light nodes must trust a centralized server (typically a full node) for submitting and confirming transactions, and checking validity of blocks.
As the Ethereum network scales, there is an increasing need support light nodes without requiring trust on a central party. Efforts such as the Portal network (Docs, 2024b) hint at the community’s desire to progress in this direction. We envision a design in which light nodes connect with other light nodes to form a gossip network. Full nodes supply block headers to light clients which gossip the header to other clients over the gossip network. In case a block is invalid, full nodes construct a fraud proof which is again gossiped over the light node network. However, a full node can construct a fraud proof only if all the data in the invalid block is available in the network. Thus, it is important for light nodes to independently verify data availability of a block lest they think the block is valid sans fraud proof.
2.2. Ethereum Peer-to-Peer Network
Ethereum uses a p2p publish-subscribe (pub-sub) network based on the GossipSub protocol (Vyzovitis et al., 2020) from the libp2p framework (Libp2p, 2024) for disseminating messages. In a pub-sub network, each node subscribes to one or more topics the node is interested in. Any message published by a node in the network belongs to a unique topic. The network is set up such that when a message is published from a topic, all nodes subscribed to that topic receive the message. In GossipSub this is achieved by constructing one independent overlay for each topic, containing all nodes subscribed to that topic. Therefore, whenever a message is published from a topic, by simply broadcasting the message over the topic’s overlay, all subscribers are guaranteed to receive the message. The primary use of a pub-sub network in Ethereum is for aggregating attestations from 64 different subsets of validators (of Stake Consensus Specifications, 2024).
In the context of data sharding, once again Ethereum plans to use a pub-sub network for data distribution and retrieval. Specifically, if a block is divided into $m$ chunks (after erasure coding), there will be $m$ topics with the $i$-th topic including the $i$-th chunk of each block. Each node subscribes to a small number of topics and is responsible for storing chunks of those topics. Chunk dissemination is straightforward: the block publisher disseminates the $i$-th chunk of the block over the $i$-th subnet. Details on how a node looking to randomly sample chunks can contact nodes from a different topic overlay to fetch chunks are unfortunately lacking in publicly available documents, forums and blog-posts.
@@ 3.1. Network and Security Model
We consider a network comprising of $n$ nodes $V$, out of which a fraction $f$ of the nodes are adversarial. Nodes that are not adversarial are called honest. We denote the set of honest and adversarial nodes by $V_{h}$ and $V_{a}$ respectively. Each node has a unique network address (i.e., IP address, port). A node can connect and communicate with another node if it knows the latter’s network address. Each node has a small memory of size $k$, for storing information about $k$ other nodes in the network. [^3] We refer to the memory space also as an address table. Apart from the network address, we assume a node also has a public, private key pair which can be used for signing messages, issuing commitments etc. Since the memory space is small, an honest node cannot know the network addresses of the entire network. However, adversarial nodes can pool together the addresses they know of to conduct attacks.
@@
When a node first joins the network, it contacts a bootstrapping server from which it receives information (network address, public key) about $k$ random nodes in the network. It is common for practical p2p networks (including Ethereum’s) to use bootstrapping servers with hardcoded IP address to help new peers join the network (Ethereum.org, 2024d). Once a node joins the network, it must run its own discovery protocol and cannot query the bootstrapping server for fresh addresses. Note that the initial set of addresses the bootstrapping server provide can include adversarial nodes as well.
We assume the public key to network address binding of a node is attested by the bootstrapping server through a certificate signed by the server. While it is possible to use a decentralized public-key infrastructure for this purpose (Shi et al., 2022), we consider the bootstrapping server as the trusted certificate authority in our setting for simplicity.
Time is divided into rounds $t=0,1,2,\ldots$. During a round, a node can send or receive messages of total size $l$, where $l$ is again a small constant that models the bandwidth constraints of the node. Lastly, we assume nodes have access to a fresh public random number each round. In practice, a new block is produced in Ethereum every 12 seconds. Thinking of 12 seconds as a round, a random number can be derived each round from the header of the block for that round. Note that we require nodes to only download the block header to compute the randomness, and not the entire block, which is particularly useful for light nodes.
@@ 3.2. Problem Statement
For any node $v\in V$, let $M_{v}(t)$ denote the contents of the address table (i.e., memory) of node $v$ at time $t$. For any $i\in\{1,2,\ldots,k\}$, let $M_{v}^{i}(t)$ be the $i$-th address in $M_{v}(t)$. Our objective is to design a decentralized algorithm $\pi$ by which an honest node can uniformly sample nodes in the network. Specifically, for any node $v\in V_{h}$, time $t$ and index $i\in\{1,2,\ldots,k\}$, we want
@@
The requirements outlined above alone can be trivially satisfied if a node downloads random addresses from the bootstrapping server and does nothing afterwards. Therefore, we qualify our objective by additionally requiring that a node must add at least one fresh sample to its address table every $\Delta>0$ rounds, i.e.,
@@
for all $t>0$. Furthermore, for any time $t$ let $R_{v}(t)$ be the most recently added address to $v$’s address table. We want the newly sampled node to be independent of the past samples, i.e.,

### v4

# Honeybee: Byzantine Tolerant Decentralized Peer Sampling with Verifiable Random Walks

@@ Abstract.
Popular blockchains today have hundreds of thousands of nodes and need to be able to support sophisticated scaling solutions—such as sharding, data availability sampling, and layer-2 methods. Designing secure and efficient peer-to-peer (p2p) networking protocols at these scales to support the tight demands of the upper layer crypto-economic primitives is a highly non-trivial endeavor. We identify decentralized, uniform random sampling of nodes as a fundamental capability necessary for building robust p2p networks in emerging blockchain networks. Sampling algorithms used in practice today (primarily for address discovery) rely on either distributed hash tables (e.g., Kademlia) or sharing addresses with neighbors (e.g., GossipSub), and are not secure in a Sybil setting. We present Honeybee, a decentralized algorithm for sampling nodes that uses verifiable random walks and table consistency checks. Honeybee is secure against attacks even in the presence of an overwhelming number of Byzantine nodes (e.g., $\geq 50$% of the network). We evaluate Honeybee through experiments and show that the quality of sampling achieved by Honeybee is significantly better compared to the state-of-the-art. Our proposed algorithm has implications for network design in both full nodes and light nodes.
blockchain, peer-to-peer, node sampling
copyright: acmlicensedpublicationmonth: 02journalyear: 2024doi: arXivconference: arXiv Preprint; 2024; 2024isbn: ccs: Networks Network algorithmsccs: Theory of computation Distributed algorithmsccs: Networks Peer-to-peer protocolsccs: Networks Network performance evaluationccs: Networks Network propertiesccs: Networks Peer-to-peer networksccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Blockchains are steadily maturing into ecosystems that support decentralized applications (dapp) in diverse domains—finance, payments, storage, games, healthcare etc.—and used by millions of clients for conducting billions of dollars of transactions each day (CoinMarketCap, 2024). As demand for dapp usage increases, it is important that the blockchains can handle the high rate of transaction requests. Today major blockchains such as Bitcoin and Ethereum can process at most a few ten transactions per second on average which has resulted in unacceptably high transaction fees during periods of high demand (Etherscan, 2024b).
To improve the scalability of blockchains, solutions such as state sharding, data sharding, and layer-2 methods (e.g., rollups) have been proposed. As these solutions are integrated into mainstream blockchains (e.g., Ethereum, Cardano, etc.), we are also seeing a big surge in the scale of the blockchain networks. For instance, today the Ethereum blockchain contains close to a million validator nodes. At these scales, designing secure and effective peer-to-peer networking protocols to support the complex upper-layer crypto-economic primitives is a significant challenge. Blockchains today are overtly reliant on random graph topologies, distributed hash tables (notably Kademlia), and p2p publish-subscribe protocols (such as GossipSub) for their networking needs. However, even the most basic problem of how to design a p2p network that eliminates eclipse attacks for a proof-of-work or proof-of-stake chain has received little attention in the literature. While heuristic approaches have offered an acceptable network performance historically, a growing network scale together with complex messaging primitives (beyond just broadcast) creates an urgent need to research principled approaches for p2p network design.
In particular, we consider the ability of a node to sample peers uniformly at random from the network without reliance on centralized entities—especially when the network contains Byzantine or Sybil nodes—as a crucial capability necessary for emerging large-scale blockchain networks. A decentralized peer-sampling service provides several benefits. It enables peers to construct (and maintain) a robust overlay that is secure against eclipsing and network partitioning attacks. Even under a high fraction of dishonest nodes (e.g., 50%), by randomly sampling and connecting to peers, honest nodes can ensure they form a connected subgraph. A tightly connected subgraph of honest nodes helps provide strong safety and liveness guarantees for the consensus algorithm running on the network. E.g., a “balancing attack” in proof-of-stake Ethereum becomes possible when adversaries form connections such that honest nodes have a lower network latency to adversarial nodes than to other honest nodes (Neuder et al., 2021; Schwarz-Schilling et al., 2022; Neu et al., 2022). This allows attackers to carefully time and release votes and blocks to honest nodes slowing down consensus. A balancing attack is also possible if the adversary can obtain accurate statistics on the message propagation latencies in the network. If peers continuously obtain fresh address samples and regularly update the overlay, such a strategy will not work.
@@
A random peer sampling service also helps with sharding and data availability sampling. E.g., to address the scaling problem, the Ethereum community is focusing on developing a data-sharding design in which (1) block sizes are increased from the current 80 kB (average) to as big as 30 MB (Nikolaenko and Boneh, 2024) and erasure coded, (2) each validator node stores only a small chunk of each coded block. Unlike Ethereum 1.0 which used a simple broadcast gossip primitive to disseminate transactions and blocks in the network or Ethereum 2.0 which uses a publish-subscribe model to disseminate different types of messages (e.g., attestations, blocks, transactions) over different subnets, a node in a data-sharded Ethereum must additionally be able to check the availability of a full block by sampling random chunks from other nodes (Al-Bassam et al., 2018). This requires a requesting node to be able to contact randomly selected nodes in the network and download chunks. Having the ability to sample random chunks is important not only for full functional nodes, but also for light nodes (e.g., a wallet running on a smart phone) which can number in the millions compared to the thousands of full nodes available today (Ethernodes.org, 2024). Peer sampling also helps blockchains employing full sharding. During a shard reconfiguration event, a migrating node must quickly discover nodes and download the latest state from the destination shard to minimize disruptions in the consensus. Last, peer sampling can help with server discovery in large-scale decentralized applications. E.g., in Livepeer a user selects a media server for encoding and distributing their video streams. A capability to select servers randomly not only helps with load balancing but can also add to security. The emerging trend of modular blockchains advocates for a separate data availability layer in the blockchain stack, the implementation of which can make use of peer sampling services as well (Celestia, 2024; EigenLayer Docs, 2024).
@@
In an open, permissionless and decentralized setting that blockchains operate on, developing an algorithm by which a node can uniformly sample other nodes in the p2p network is far from straightforward. At present, the two common approaches for peer discovery are using (1) the Kademlia distributed hash table (DHT) protocol (Maymounkov and Mazieres, 2002), or (2) p2p exchange of addresses between neighbors in the overlay (e.g., used by GossipSub). To discover a random node in the network with Kademlia, all a node has to do is issue a query for a randomly selected target identifier and receive the IP address information of the node that is closest to the target identifier (in Kademlia’s XOR distance sense) (Ethereum Peer-to-Peer Networking Specifications, 2024b, a).
@@
In this paper we argue that in the presence of adversarial nodes (or Sybils), neither of the currently followed approaches achieve uniform sampling and can cause a requesting node to become eclipsed. As a potential replacement, in this paper we propose Honeybee which is a fully decentralized p2p algorithm for performing uniform node sampling even in the presence of adversarial nodes. A Honeybee node achieves sampling by participating in several random walks over the p2p overlay. Each node in Honeybee maintains an address table containing addresses of the most recently sampled peers which are also used to progress random walks of other peers visiting the node. To protect against adversarial attacks, Honeybee uses a verifiable public source of randomness (e.g., can be derived from the blockchain) to perform the walks. Additionally, Honeybee nodes also perform peer-to-peer reconciliation of node address tables to identify and expose attackers engaging in equivocating their address tables.
@@
We do not attempt to present a full-fledged p2p network design for the applications outlined above (broadcast, data availability sampling, sharding, server discovery etc.) in this paper. Rather we posit that a uniform node sampling capability will have a central role to play in the eventual overall network design for these applications.
@@
We evaluate Honeybee through a custom simulator we have built. [^1] Code will be made open source. Compared to the baseline algorithms GossipSub and Kademlia, Honeybee achieves the same level of near-uniform sampling when all nodes in the network adhere to the protocols. However, when the network contains adversarial nodes, Honeybee outperforms GossipSub and Kademlia by 4-63% in terms of sampling adversarial peer ratio. We define that an algorithm achieves $\epsilon$-uniform sampling when the sampling adversarial peer ratio from the algorithm is bounded from above by the sum of the true adversarial nodes ratio in the network and $\epsilon$. Under such standard, Honeybee consistently achieves 0.03-uniform sampling with 5-70% adversarial nodes in the network.
@@ 2.1. Network and Security Model
We consider a network comprising of $n$ nodes, out of which a fraction $f$ of the nodes are adversarial (dishonest). Nodes that are not adversarial are called honest. We denote the set of honest and adversarial nodes by $V_{h}$ and $V_{a}$ respectively. $V$ denotes the set of all nodes. Each node has a unique network address (i.e., IP address, port). A node can connect and communicate with another node if it knows the latter’s network address. Each node has a small memory of size $k$, for storing information about $k$ other nodes in the network. [^2] We refer to the memory space also as an address table. Apart from the network address, we assume a node also has a public, private key pair used for signing messages and issuing commitments. Since the memory space is small, an honest node cannot know the network addresses of the entire network. However, adversarial nodes can pool together the addresses they know of to conduct attacks.
@@
When a node first joins the network, it contacts a bootstrapping server from which it receives information (network address, public key) about $k$ random nodes in the network. It is common for practical p2p networks (e.g., Ethereum’s network) to use bootstrapping servers with hardcoded IP address to help new peers join the network (Ethereum.org, 2024b). Once a node joins the network, it must run its own discovery protocol and cannot query the bootstrapping server for fresh addresses. Note that the initial set of addresses the bootstrapping server provides can include adversarial nodes as well.
We assume each IP address has an associated unique public key. The public key to IP address binding of a node is attested by the bootstrapping server through a certificate signed by the server. While it is possible to use a decentralized public-key infrastructure for this purpose (Shi et al., 2022), we consider the bootstrapping servers as the trusted certificate authority in our setting for simplicity.
Time is divided into rounds $t=0,1,2,\ldots$. During a round, a node can send or receive messages of total size $l$, where $l$ is again a small constant that models the bandwidth constraints of the node. Lastly, we assume nodes have access to a fresh public random number each round. In practice, a new block is produced in Ethereum every 12 seconds. Thinking of 12 seconds as a round, a random number can be derived each round from the header of the block for that round. Note that we require nodes to only download the block header to compute the randomness, and not the entire block, which is particularly useful for light nodes. Alternative public and trusted sources of randomness may also be used as we will discuss in §4.3.
@@ 2.2. Problem Statement
For any node $v\in V$, let $M_{v}(t)$ denote the contents of the address table (i.e., memory) of node $v$ at time $t$. For any $i\in\{1,2,\ldots,k\}$, let $M_{v}^{i}(t)$ be the $i$-th address in $M_{v}(t)$. Our objective is to design a decentralized algorithm by which an honest node can continuously sample nodes uniformly at random from the network. Specifically, for any node $v\in V_{h}$, time $t$ and index $i\in\{1,2,\ldots,k\}$, we want
@@
The requirements outlined above can be trivially satisfied if a node downloads random addresses from the bootstrapping servers once initially and does not update the addresses afterward. Therefore, we qualify our objective by additionally requiring that a node must add at least one fresh sample to its address table every $\Delta>0$ rounds, i.e.,
@@
for all $t>0$. In addition, since it is infeasible for the bootstrapping servers to handle the overhead as network size increases, we preclude the trivial solution of refreshing addresses by periodically downloading random addresses from the bootstrapping servers. Furthermore, for any time $t$, let $R_{v}(t)$ be the most recently added address to $v$’s address table. We want the newly sampled node to be independent of the past samples, i.e.,